OpenStudy (anonymous):

PLS HELP! MEDALS AND FAN! I'm Dying! Use elimination to find the solution to the system of equations. 4x+y=8 3x-2y=6 A. x=5,y=-12 B. x= 16/3, y=5 C. x=2, y=0 D. x=-2, y=16

OpenStudy (anonymous):

(4x+y=8)2 8x+2y=16 3x-2y=6 add the equations 11x=22 x=2 y is being cancelled substitute x 4x+y=8 4(2)+y=8 y=0 x=2, y=0

OpenStudy (anonymous):

Thank you!